# Multitype entanglement dynamics induced by exceptional points

### v1

@@ IV Different entanglement dynamics
$\frac{d\rho}{dt}=-i\left(H\rho-\rho{{H}^{\dagger}}\right)$ (3)
where $H$ is the non-Hermitian system Hamiltonian [Eq. (1)] and $\rho$ denotes the density matrix of the system. To describe the entanglement between the two NH qubits, we calculate the concurrence as Wootters (1998)
@@
Here ${{\tau}_{1}}$, ${{\tau}_{2}}$, ${{\tau}_{3}}$, ${{\tau}_{4}}$ are the eigenvalues of the Hermitian matrix $R=\sqrt{\sqrt{\rho}\widetilde{\rho}\sqrt{\rho}}$, with $\widetilde{\rho}=\left({{\sigma}_{y}}\otimes{{\sigma}_{y}}\right){{\rho}^{*}}\left({{\sigma}_{y}}\otimes{{\sigma}_{y}}\right)$ in decreasing order and ${{\sigma}_{y}}=-i\left|b\right\rangle\left\langle a\right|+i\left|a\right\rangle\left\langle b\right|$.
Figure 3: The concurrence evolution in the region of PTB (a), the new phase (b) and PTS (c) in the case of weak coupling $\xi/\gamma=0.006$. The original EP4 induce the exceptional entanglement phenomena on the two sides of this point. In the strong coupling regime for $\xi/\gamma=0.5$, the EP2 induces different entanglement behavior on the two sides of $\Omega/\gamma=0.626$, i.e., in the region of PTS (d, f) and PTB (e). The initial state is $\left|\psi(0)\right\rangle=\left|aa\right\rangle$.
Assuming the initial state of the two coupled NH qubits is $\left|\psi(0)\right\rangle=\left|aa\right\rangle$, the concurrence evolution in both the weak [Fig. 3(a),  3(b) and  3(c)] and strong coupling regime [Fig. 3(d),  3(e) and  3(f)] can be obtained numerically. The entanglement behaviors are significantly different when tuning different parameters. In the case of weak coupling (taking $\xi/\gamma=0.006$ as an example), the degree of entanglement is weak and monotonically increases until evolving to a stable value in the region of $\Omega<{{\Omega}_{\text{EP4}^{\prime}}}$ [Fig. 3(a)]. We define this type of entanglement dynamics as type I, which can only be observed in the weak coupling regime. With increasing the driving amplitude to the second region, i.e., ${{\Omega}_{\text{EP4}^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$, the entanglement dynamics exhibits a combination function of decay and oscillation, reaching to a fixed value of concurrence at the steady state [Fig. 3(b)]. Clearly, this type of entanglement dynamics is different from the type I, which we defined as type III.
According to the results in Fig. 3(a) and  3(b), we can find that the entanglement behaviors are absolutely different when crossing the orginal EP4, which indicates that the entanglement dynamics transition is induced by the original EP4, although the original EP4 is absent in the presence of coupling. It is necessary to define this particular region where enclosed by the original EP4 and EP2 as a new phase, as the blue region shown in Fig. 1(b). This new phase is a part of the PTB.
In the region of PTS, i.e., $\Omega>{{\Omega}_{\text{EP2}}}$, the entanglement dynamics exhibits continuous oscillation in the whole time-evolution process [Fig. 3(c)], which is defined as type II. The maximum entanglement ($C=1$) can be obtained in the oscillation process and the degree of entanglement in PTS is much larger than that in PTB. By using the theory of time-independent perturbation theory for non-Hermitian systems Li et al. (2023b), we can obtain the analytical results of the concurrence in the weak coupling regime, which shows a good agreement with the numerical results in the $\mathcal{P}\mathcal{T}$-symmetric phase (see Appendix C.3 for more details). The multitype entanglement dynamics (type I, II, III) as described above can also be seen in the aspect of the population of the basis state. The normalized quantum state can be expressed as $\left|\widetilde{\psi}\right\rangle=\alpha\left|aa\right\rangle+\beta\left|ab\right\rangle+\zeta\left|ba\right\rangle+\delta\left|bb\right\rangle$. Hence the concurrence is given by $C=2\left|\alpha\delta-\beta\zeta\right|$. corresponding to the population at each basis state exhibits different oscillatory behaviors in the three phases divided by the original EP4 and EP2 (as shown in Fig. 4), different entanglement dynamics can be found in the three phases, accordingly.
Figure 4: Time evolution of the modulus $\left|j\right|(j=\alpha,\beta,\varsigma,\delta)$ for each of the four complex amplitudes of the two-qubit state $\left|\widetilde{\psi}\right\rangle=\alpha\left|aa\right\rangle+\beta\left|ab\right\rangle+\zeta\left|ba\right\rangle+\delta\left|bb\right\rangle$ for $\xi=0$ (a) and $\xi/\gamma=0.006$ (b-d). In the absence of Ising interaction, the evolution of $\left|j\right|$ is focused in the $\mathcal{P}\mathcal{T}$-symmetric phase. In the presence of Ising interaction, the evolution of $\left|j\right|$ has been shown in three different regions, i.e., in the $\mathcal{P}\mathcal{T}$-symmetry-broken phase (b), the new phase (c) and $\mathcal{P}\mathcal{T}$-symmetric phase (d), respectively. The given driving amplitude are chosen as $\Omega/\gamma=0.23$ (b), $\Omega/\gamma=0.28$ (c) and $\Omega/\gamma=0.3$ (d).
Compared with the the populations evolution of the basis states $\left|bb\right\rangle$ and $\left|aa\right\rangle$ in the three different phases, we can find that the distorted Rabi-like oscillation of the population of $\left|bb\right\rangle$ and $\left|aa\right\rangle$ in the PTS, as the black and pink curves shown in Fig. 4(a) and  4(d) regardless of whether there exists interaction. Moreover, the population evolution of the states $\left|ab\right\rangle$ and $\left|ba\right\rangle$ monotonically increase during the system evolving to a steady state in the PTB [Fig. 4(b)] and the new phase [Fig. 4(c)]. Strikingly, the populations evolution of the basis states $\left|bb\right\rangle$ and $\left|aa\right\rangle$ are oscillations at short timescale and finally reach to a steady state in the new phase, which are different from that in the PTB. The different populations evolution corresponds to the distinctive entanglement dynamics in the region of new and PTB. In addition, we note that at specific time ${{T}^{*}}=27.2$, the populations of basis state are equal, i.e., $\left|\alpha\right|\sim\left|\beta\right|\sim\left|\varsigma\right|\sim\left|\delta\right|$.
@@
With the increase of the coupling strength to the strong coupling regime, i.e., $\xi/\gamma=0.5$ as an example, the original EP4 is replaced by another EP2, which means there are two different EP2s in the system [see Fig. 2(c) and  2(f)]. The entanglement between two coupled NH qubits has been shown in Fig. 3(d),  3(e) and  3(f). We can observe that the entanglement dynamics of type II takes place both in the region of $\Omega<\Omega_{\text{EP2}}^{(1)}$ and $\Omega>\Omega_{\text{EP2}}^{(2)}$, i.e., the region of PTS, as shown in Fig. 3(d) and  3(f), respectively. In PTB, i.e., $\Omega_{\text{EP2}}^{(1)}<\Omega<\Omega_{\text{EP2}}^{(2)}$, the entanglement dynamics is the same as the defined type III as demonstrated in Fig. 3(e). Therefore, only two types of entanglement dynamics can be found in the strong coupling regime. When calculating the entanglement dynamics in the regions near the two sides of the original EP4 ($\Omega/\gamma=0.25$), which behaves almost consistent (see Fig. 5(a) and  5(b)). It means that due to the replacement of the original EP4 with another EP2, entanglement dynamics of type I will not appear in the strong coupling regime. The impact of original EP4 on entanglement disappears, which also indicates that the new phase vanishes.
Figure 5: Concurrence evolution for the critical values near the original EP4 in the strong coupling regime. As an example, for $\Omega/\gamma=0.23$ (a) and $\Omega/\gamma=0.28$ (b) in the left and right sides near the original EP4, respectively.
The effect of the original EP4 on entanglement dynamics are completely different in the weak and strong coupling regime. The presence of the weak coupling can be considered as a perturbation, such that the original EP4 still makes a significant influence on the parameter space, leading to a different entanglement dynamics on the two sides near the original EP4. However, when the coupling strength is large enough, the contribution of the strong coupling on the parameter space is dominant, leading to the disappearance of the entanglement dynamics transition induced by the original EP4. Therefore, the coexistence of the two EPs with different orders, i.e., EP4 and EP2, can induce three different types of entanglement dynamics in the weak coupling regime. Depend on the same order of the two EP2s, only a jump of entanglement dynamics can be found in the strong coupling regime.
V Entanglement transition without the driving field
@@
In the absence of the driving field applied to the qubits, i.e., ${{\Omega}_{i=1,2}}=0$. Exceptional entanglement transition can also occur when considering the Ising type interaction. The Hamiltonian with Ising type interaction evolves within the subspace $\left\{\left|{{a}_{1}},{{a}_{2}}\right\rangle,\left|{{a}_{1}},{{b}_{2}}\right\rangle,\left|{{a}_{2}},{{b}_{1}}\right\rangle,\left|{{b}_{1}},{{b}_{2}}\right\rangle\right\}$. With a correction of the quantum state distortion caused by the decoherence, the evolution of the joint probability $\left|{{b}_{1}},{{b}_{2}}\right\rangle$, denoted as ${{P}_{{{b}_{1}}{{b}_{2}}}}$, indicating the quantum Rabi oscillator signal. Figure 6(a) and  6(b) show the obtained ${{P}_{{{b}_{1}}{{b}_{2}}}}$ as functions of time in the regions of different phases. Specifically, in PTB (e.g., $\xi/\gamma=-0.16$), the population of state $\left|{{b}_{1}},{{b}_{2}}\right\rangle$ monotonically decays during the system evolving to a steady state [Fig. 6(a)]. After crossing the EP, i.e., in the region of PTS (e.g., $\xi/\gamma=-0.5$), the state population ${{P}_{{{b}_{1}}{{b}_{2}}}}$ presents a distinct oscillatory behavior in time evolution as shown in Fig. 6(b).
@@
Figure 6: Observation of exceptional phase transitions in the absence of $\Omega$. (a) and (b) Measured evolution of the population of the basis state $\left|{{b}_{1}},{{b}_{2}}\right\rangle$ in PTB (for $\xi/\gamma=-0.16$) and PTS (for $\xi/\gamma=-0.5$), respectively. Concurrence evolution are given for $\xi=-0.2\gamma$ (c) and $\xi=-0.5\gamma$ (d). (e) Concurrence values ${{\varepsilon}_{\pm}}$ for the eigenstates $\left|{{\Phi}_{+}}\right\rangle$ and $\left|{{\Phi}_{-}}\right\rangle$ versus the Ising type interaction strength $\xi$. The derivatives $d{{\varepsilon}_{\pm}}/d\xi$ around the EP2, obtained by $\left[{{\varepsilon}_{\pm}}(\xi+\delta\xi)-{{\varepsilon}_{\pm}}(\xi)\right]/d\xi$, as shown in the insets. (f) Spectral gap $\Delta{{E}_{g}}$. As the eigenspectrum is possessed the entangled eigenstates of the two qubits system. This gap corresponds to the vacuum Rabi splitting.
@@
To show the relationship between entanglement behavior and the EP, we give the concurrence evolution for $\xi=-0.2\gamma$ and $-0.5\gamma$ in Fig. 6(c) and  6(d), respectively. The result demonstrates that the entanglement exhibits different evolution behaviors in PTB and PTS. Figure 6(e) indicates the concurrence values ${{\varepsilon}_{\pm}}$ associated with the two eigenstates $\left|{{\Phi}_{\pm}}\right\rangle$ of the system as functions of $\xi$. As theoretical predicted, below the EP (i.e., $\left|\xi/\gamma\right|>{\gamma}/{4}\;$), the concurrence value of each eigenstate is saturated and approximately converge to the same maximally entangled state (reaching the maximum value 1) and independent of $\xi$ until at the EP $\xi=-\gamma/4$, where the energy gap disappeared. The expression of the concurrence has been given in Appendix A. The discontinuity of these derivatives indicates the occurrence of an entanglement transition at the EP, which are shown in the insets. After crossing the EP (i.e., $\left|\xi/\gamma\right|<{\gamma}/{4}\;$), the concurrence values ${{\varepsilon}_{\pm}}$ exhibit a linear scaling and depends on $\xi$. ${{\varepsilon}_{\pm}}$ are decreasing to 0 with the reduce of $\left|\xi\right|$. In Fig. 6(f), the energy gap described as ${{E}_{g}}=\sqrt{4{{J}^{2}}-{{\gamma}^{2}}}$ with $J=-2\xi$, undergoes a transition from the real part to the imaginary part at the EP, which is accompanied by an entanglement transition of the eigenstates. In addition, according to the results in Fig. 6(a), 6(b) and  6(e), this type of the energy gap transition also accompanied by the effect of the vacuum Rabi splitting of the entangled states.
VI The calculation for the system with dipolar interaction

### v2

@@ IV Different entanglement dynamics
$\frac{d\rho}{dt}=-i\left({{H}_{j=1,2}}\rho-\rho H_{j=1,2}^{\dagger}\right),$ (3)
where $\rho$ denotes the density matrix of the system. Concurrence is a useful tool for quantifying the entanglement of bipartite quantum systems, which can be written as Wootters (1998)
@@
Here ${{\tau}_{1}}$, ${{\tau}_{2}}$, ${{\tau}_{3}}$, ${{\tau}_{4}}$ are the eigenvalues of the Hermitian matrix $R=\sqrt{\sqrt{\rho}\widetilde{\rho}\sqrt{\rho}}$, with $\widetilde{\rho}=\left({{\sigma}_{y}}\otimes{{\sigma}_{y}}\right){{\rho}^{*}}\left({{\sigma}_{y}}\otimes{{\sigma}_{y}}\right)$ in decreasing order, ${{\rho}^{*}}$ is the complex conjugation of the density matrix $\rho$ and ${{\sigma}_{y}}=-i\left|b\right\rangle\left\langle a\right|+i\left|a\right\rangle\left\langle b\right|$.
Figure 3: As an example for weak coupling strength $\xi/\gamma=0.006$, the concurrence evolution in the region of PI (a), the mixed phase (b) and PTS (c). The original EP4 and EP2 induce the entanglement dynamics change on the two sides of these points. In the strong coupling regime for $\xi/\gamma=0.5$, the EP2 induces the spectrum transition of the PTS and mixed phase, leading to the type II (d, f) and III (e) entanglement dynamics change. All plots are made for $\left|\psi(0)\right\rangle=\left|aa\right\rangle$. The dashed curves denote the envelope fitting of the concurrence evolutions.
For a given initial state $\left|\psi(0)\right\rangle$, the normalized quantum state can be expressed as $\left|\widetilde{\psi}(t)\right\rangle=\alpha(t)\left|aa\right\rangle+\beta(t)\left|ab\right\rangle+\zeta(t)\left|ba\right\rangle+\delta(t)\left|bb\right\rangle$. Hence, the concurrence is given by
$C=2\left|\alpha\delta-\beta\zeta\right|.$ (5)
Hence, the concurrence can be used to calculate the degree of entanglement between the qubits (more details are given in Appendix C.3).
By solving Eq. (3) and Eq. (4), the numerical results of the concurrence evolution are displayed in Fig. 3. In the case of weak coupling (taking $\xi/\gamma=0.006$ as an example), the entanglement dynamics show exponential-like behavior in the region of $\Omega<{{\Omega}_{\text{EP4}^{\prime}}}$ [Fig. 3(a)]. We define this type of entanglement dynamics as type I, which can only be observed in the weak coupling regime. In the region of PTS, i.e., $\Omega>{{\Omega}_{\text{EP2}}}$, the entanglement dynamics exhibits continuous oscillation in the whole time-evolution process [Fig. 3(c)], which is defined as type II. By fitting envelope functions to the type II, the most obvious characteristic of this type entanglement is that both the upper and lower branches exhibit significant oscillatory behavior. In the mixed phase, the feature of the entanglement dynamics exhibits a combination of both the exponential-like and oscillatory behavior [Fig. 3(b)]. Clearly, this type of entanglement dynamics is defined as type III. One can distinguish the type III entanglement dynamics in the strong coupling and weak coupling regimes by comparing the decay rate of the upper and lower envelopes (see Fig. C3).
Meanwhile, the concurrence evolution is ranging from 0 to 1, and the maximum entanglement ($C=1$) in PTS is much larger than that in PI and the mixed phase. By using the theory of time-independent perturbation theory for non-Hermitian systems Z.-Z. Li, W. Chen, M. Abbasi, K. W. Murch, and K. B. Whaley (2023), we can obtain the analytical results of the concurrence in the weak coupling regime, which shows a good agreement with the numerical results in the $\mathcal{P}\mathcal{T}$-symmetric phase (see Appendix C.3 for more details). The different features of the entanglement dynamics also provides an efficient approach to distinguish the spectrum phase (PTS, PTB) of the system.
Furthermore, we display that on approaching the EP4’ in PI (i.e., $\Omega<{{\Omega}_{\text{EP}4^{\prime}}}$), the speed of reaching steady-state entanglement decreases, meaning that the required time becomes longer, i.e., ${{T}_{1}}<{{T}_{2}}<{{T}_{3}}$ (see Fig. 4(a)). On the contrary, in the mixed phase (i.e., ${{\Omega}_{\text{EP}4^{\prime}}}<\Omega<{{\Omega}_{\text{EP2}}}$), when forcing proximity to the EP4’, steady-state entanglement can be achieved more quickly (see Fig. 4(b)). Based on this observation, although both the PI and the mixed phase fall within the $\mathcal{P}\mathcal{T}$-symmetry-broken phase, the steady-state entanglement generation near higher-order EPs behaves in completely opposite ways in these two regions.
Figure 4: The concurrence evolution in the PI (a) and the mixed phase (b). In PI, the observation of steady-state entanglement generation at a longer timescale by proximity to the original EP4, as the black arrow shown. However, in the mixed phase, the steady-state entanglement can be generated at a speed much faster when approach to the original EP4. The other parameter is chosen as $\xi/\gamma=0.006$ and the initial state is $\left|\psi(0)\right\rangle=\left|aa\right\rangle$.
This phenomenon can be understood as follows. On the one hand, if the applied driving amplitude smaller than the original EP4, i.e., $\Omega<{{\Omega}_{\text{EP}4^{\prime}}}$, even a smaller driving amplitude can lead to realize faster steady-state entanglement generation. On the other hand, once the driving amplitude larger than the original EP4, i.e., $\Omega>{{\Omega}_{\text{EP}4^{\prime}}}$, where the EP4’ can be treated as a perturbation in the parameter space. The greater the driving amplitude, the larger the oscillations in the concurrence evolution, resulting in a longer timescale to achieve steady-state entanglement.
@@
With the increase of the coupling strength to the strong coupling regime, i.e., $\xi/\gamma=0.5$ as an example, the original EP4 is replaced by another EP2, which means there are two different EP2s in the system [see Figs. 2(c) and  2(f)]. The entanglement between two coupled NH qubits has been shown in Figs. 3(d)  3(f). We can observe that the entanglement dynamics of type II takes place in the region of $\Omega<\Omega_{\text{EP2}}^{(1)}$ and $\Omega>\Omega_{\text{EP2}}^{(2)}$, i.e., the region of PTS, as shown in Figs. 3(d) and  3(f), respectively. In the region of mixed phase, i.e., $\Omega_{\text{EP2}}^{(1)}<\Omega<\Omega_{\text{EP2}}^{(2)}$, the type III entanglement dynamics can be observed, as shown in Fig. 3(e). Therefore, only two types of entanglement dynamics can be found in the strong coupling regime.
V The discussion of the entanglement dynamics change
The normalized quantum state evolved under the system Hamiltonian Eq. (1) in the main text can be written as
$\left|\widetilde{\psi}(t)\right\rangle=\sum\limits_{n}{{{c}_{n}}{{e}^{-i{{\lambda}_{n}}t}}}\left|{{\lambda}_{n}}\right\rangle and{{\lambda}_{n}}=\operatorname{Re}({{\lambda}_{n}})+i\operatorname{Im}({{\lambda}_{n}}),$ (6)
where ${{\lambda}_{n}}$ and $\left|{{\lambda}_{n}}\right\rangle$ are eigenvalues and eigenstates of the non-Hermitian Hamiltonian Eq. (1), respectively. Given an initial state $\left|\psi(0)\right\rangle$, the time evolution of the state under the system Hamiltonian $U={{e}^{-iHt}}$ can be read as
$\left|\widetilde{\psi}(t)\right\rangle=U\left|\psi(0)\right\rangle=\sum\limits_{n=1}^{4}{\left\langle{{\lambda}_{n}}|\psi(0)\right\rangle\cdot{{e}^{-it{{\lambda}_{n}}}}\left|{{\lambda}_{n}}\right\rangle},$ (7)
Correspondingly, we can obtain the coefficient of the state as $\alpha(t)={{c}_{1}}\cdot{{e}^{-it{{\lambda}_{1}}}},\beta(t)={{c}_{2}}\cdot{{e}^{-it{{\lambda}_{2}}}},\varsigma(t)={{c}_{3}}\cdot{{e}^{-it{{\lambda}_{3}}}},\delta(t)={{c}_{4}}\cdot{{e}^{-it{{\lambda}_{4}}}}$, with the coefficients determined by the initial state as ${{c}_{1}}=\left\langle{{\lambda}_{1}}|\psi(0)\right\rangle$, ${{c}_{2}}=\left\langle{{\lambda}_{2}}|\psi(0)\right\rangle$, ${{c}_{3}}=\left\langle{{\lambda}_{3}}|\psi(0)\right\rangle$ and ${{c}_{4}}=\left\langle{{\lambda}_{4}}|\psi(0)\right\rangle$. In this case, the concurrence evolution can be given by
$C(t)$ $=2\left|c_{1}c_{4}e^{-it(\lambda_{1}+\lambda_{4})}-c_{2}c_{3}e^{-it(\lambda_{2}+\lambda_{3})}\right|$
$=2\left|c_{1}c_{4}e^{-it\operatorname{Re}(\lambda_{1}+\lambda_{4})}-c_{2}c_{3}e^{-it\operatorname{Re}(\lambda_{2}+\lambda_{3})}\right|$
$\quad+2\left|c_{1}c_{4}e^{t\operatorname{Im}(\lambda_{1}+\lambda_{4})}-c_{2}c_{3}e^{t\operatorname{Im}(\lambda_{2}+\lambda_{3})}\right|.$ (8)
It can be seen that the imaginary part $\operatorname{Im}({{\lambda}_{n}})$ will cause the concurrence evolution exhibiting exponential-like behavior, while the real part $\operatorname{Re}({{\lambda}_{n}})$ will cause the concurrence evolution showing oscillatory behavior.
According to Eq. (8), we can find that the eigenvalues of the system undergoes a transition at EPs from purely real to purely imaginary or complex, the evolution of the concurrence also changes. Specifically, we compare the different characteristic of eigenvalues in three regions (i.e., the PI, the mixed phase and the PTS). The original EP4 induces the spectrum transition of the PI and mixed phase, leading to the type I (exponentially-like behavior, see Fig. 3(a)) and III (combination of both the exponential-like decay and oscillatory behavior, see Fig. 3(b, e)) entanglement dynamics change in the weak coupling regime. Similarly, the EP2 induces the spectrum transition of the PTS and mixed phase, leading to the type II (oscillatory behavior, see Fig. 3(c, d, f)) and III entanglement dynamics change. This effect can be observed in both the weak and strong coupling regimes.
Figure 5: The concurrence value ${{\varepsilon}_{4}}$ in the weak coupling regime (a) and ${{\varepsilon}_{3}}$ in the strong coupling regime (b) as a function of the driving amplitude $\Omega$. The yellow circles represent the EP2 in the system while the green circles denote the original EP4.
As the eigenstates are entangled states, the concurrence about eigenstates can be treated as another manifestation of the entanglement dynamics change [see Fig.  5]. By giving concurrence associated with the fourth eigenstate ${{\varepsilon}_{4}}$ in the weak coupling regime [Fig.  5(a)] and the third eigenstate ${{\varepsilon}_{3}}$ [Fig.  5(b)] in the strong coupling regime, we demonstrate that there is a notable inflection point appears in the concurrence evolution at the EP2. In particular, the slope of the concurrence evolution changes abruptly at the inflection point (i.e., EP2), suggesting that the occurrence of an entanglement dynamics change at the EP2. In addition, we note that there is a notable change in the trend of the concurrence evolution near the EP4’ in both weak and strong coupling regimes, as the red and blue arrows shown. As a result, the entanglement dynamics change between type I and II can be observed at the EP4’.
VI Entanglement changing without the driving field
@@
In the absence of the driving field applied to the qubits, i.e., ${{\Omega}_{i=1,2}}=0$. Exceptional entanglement changing can also occur when considering the Ising type interaction. The Hamiltonian with Ising type interaction evolves within the subspace $\left\{\left|{{a}_{1}},{{a}_{2}}\right\rangle,\left|{{a}_{1}},{{b}_{2}}\right\rangle,\left|{{a}_{2}},{{b}_{1}}\right\rangle,\left|{{b}_{1}},{{b}_{2}}\right\rangle\right\}$. With a correction of the quantum state distortion caused by the decoherence, the evolution of the joint probability $\left|{{b}_{1}},{{b}_{2}}\right\rangle$, denoted as ${{P}_{{{b}_{1}}{{b}_{2}}}}$, indicating the quantum Rabi oscillator signal. Figure 6(a) and  6(b) show the obtained ${{P}_{{{b}_{1}}{{b}_{2}}}}$ as functions of time in the regions of different phases. Specifically, in PTB (e.g., $\xi/\gamma=-0.16$), the population of state $\left|{{b}_{1}},{{b}_{2}}\right\rangle$ monotonically decays during the system evolving to a steady state [Fig. 6(a)]. After crossing the EP, i.e., in the region of PTS (e.g., $\xi/\gamma=-0.5$), the state population ${{P}_{{{b}_{1}}{{b}_{2}}}}$ presents a distinct oscillatory behavior in time evolution as shown in Fig. 6(b).
@@
Figure 6: Observation of exceptional phase transitions in the absence of $\Omega$. (a) and (b) Measured evolution of the population of the basis state $\left|{{b}_{1}},{{b}_{2}}\right\rangle$ in PTB (for $\xi/\gamma=-0.16$) and PTS (for $\xi/\gamma=-0.5$), respectively. Concurrence evolution are given for $\xi=-0.2\gamma$ (c) and $\xi=-0.5\gamma$ (d). (e) Concurrence values ${{\varepsilon}_{\pm}}$ for the eigenstates $\left|{{\Phi}_{+}}\right\rangle$ and $\left|{{\Phi}_{-}}\right\rangle$ versus the Ising type interaction strength $\xi$. The derivatives $d{{\varepsilon}_{\pm}}/d\xi$ around the EP2, obtained by $\left[{{\varepsilon}_{\pm}}(\xi+\delta\xi)-{{\varepsilon}_{\pm}}(\xi)\right]/d\xi$, as shown in the insets. (f) Spectral gap $\Delta{{E}_{g}}$. The eigenspectrum is possessed the entangled eigenstates of the two qubits system. This gap corresponds to the vacuum Rabi splitting.
@@
To show the relationship between entanglement behavior and the EP, we give the concurrence evolution for $\xi=-0.2\gamma$ and $-0.5\gamma$ in Fig. 6(c) and  6(d), respectively. This result demonstrates that the entanglement exhibits different evolution behaviors in PTB and PTS. Figure 6(e) indicates the concurrence values ${{\varepsilon}_{\pm}}$ associated with the two eigenstates $\left|{{\Phi}_{\pm}}\right\rangle$ of the system as functions of $\xi$. As theoretical predicted, below the EP (i.e., $\left|\xi/\gamma\right|>{\gamma}/{4}\;$), the concurrence value of each eigenstate is saturated and approximately converges to the same maximally entangled state (reaching the maximum value 1) and independent of $\xi$ until at the EP $\xi=-\gamma/4$, where the energy gap disappeared. The expression of the concurrence has been given in Appendix A. The discontinuity of these derivatives indicates the occurrence of an entanglement changing at the EP, which are shown in the insets. After crossing the EP (i.e., $\left|\xi/\gamma\right|<{\gamma}/{4}\;$), the concurrence values ${{\varepsilon}_{\pm}}$ exhibit a linear scaling and depends on $\xi$. ${{\varepsilon}_{\pm}}$ are decreasing to 0 with the reduction of $\left|\xi\right|$. In Fig. 6(f), the energy gap described as ${{E}_{g}}=\sqrt{4{{J}^{2}}-{{\gamma}^{2}}}$ with $J=-2\xi$, undergoes a transition from the real part to the imaginary part at the EP, which is accompanied by an entanglement transition of the eigenstates. In addition, according to the results in Figs. 6(a), 6(b) and  6(e), this type of the energy gap transition is also accompanied by the effect of the vacuum Rabi splitting of the entangled states.
VII The calculation for the system with dipolar interaction